OpenStudy (anonymous):

let f(v) be the amount of energy consumed by a flying bird (joules/sec), as a function of its speed v (meters/sec). Let a(v) be the amount of energy consumed by the bird in (joules/meter) a) where is a(v) a minimum? let f(v) be the amount of energy consumed by a flying bird (joules/sec), as a function of its speed v (meters/sec). Let a(v) be the amount of energy consumed by the bird in (joules/meter) a) where is a(v) a minimum? @Mathematics

OpenStudy (eyust707):

f(v) = ??? J/s v = ??m/s a(v) = ??? J/m f(v)/v = (j/s)/(m/s) = j/m = a(v)

OpenStudy (eyust707):

im very very uncertian about that.. lol but i believe a(v) in smallest when v is really really big or when f(v) is really really small? honestly i would rely on this tho cause im not very sure